# HEAM : Hashed Embedding Acceleration using Processing-In-Memory

### 21 Feb 2024 v2

@@ Abstract.
In today’s data centers, personalized recommendation systems face challenges such as the need for large memory capacity and high bandwidth, especially when performing embedding operations. Previous approaches have relied on DIMM-based near-memory processing techniques or introduced 3D-stacked DRAM to address memory-bound issues and expand memory bandwidth. However, these solutions fall short when dealing with the expanding size of personalized recommendation systems. Recommendation models have grown to sizes exceeding tens of terabytes, making them challenging to run efficiently on traditional single-node inference servers. Although various algorithmic methods have been proposed to reduce embedding table capacity, they often result in increased memory access or inefficient utilization of memory resources.
@@
This paper introduces HEAM, a heterogeneous memory architecture that integrates 3D-stacked DRAM with DIMM to accelerate recommendation systems in which compositional embedding is utilized—a technique aimed at reducing the size of embedding tables. The architecture is organized into a three-tier memory hierarchy consisting of conventional DIMM, 3D-stacked DRAM with a base die-level Processing-In-Memory (PIM), and a bank group-level PIM incorporating a Look-Up Table. This setup is specifically designed to accommodate the unique aspects of compositional embedding, such as temporal locality and embedding table capacity. This design effectively reduces bank access, improves access efficiency, and enhances overall throughput, resulting in 6.3×6.3$\times$6.3 × speedup and 58.9% energy savings compared to the baseline.
1. Introduction
@@
Personalized recommendation serves as an important technology of various companies such as Meta (naumov2019deep, ), YouTube (covington2016deep, ) and Netflix (steck2021deep, ). Recently, deep learning-based recommendation systems have emerged as a pivotal technology for improving the accuracy of predicting user-preferred content, thereby contributing to enhanced profitability for these companies. These advanced recommendation systems currently contribute substantially to the computational workload of the data center. Recent reports indicate that approximately 80% of data center resources are dedicated to the inference process of recommendation systems (gupta2020architectural, ). Recommendation models use dense features for user information and sparse features for item specifics through embedding operations. Following feature interactions, multi-layer perceptron (MLP) layers generate click-through-rate (CTR) predictions, estimating the likelihood of user clicks on recommended items.
@@
In large-scale data centers that extensively utilize embeddings, the embedding operation often becomes the primary factor influencing overall system performance. This stems mainly from the operation’s tendency to exhibit sparse and irregular memory access patterns. As a result, these memory-bound operations place significant constraints on traditional data center infrastructures (zhao2022understanding, ), and this challenge continues to grow with the scale, as depicted in Figure 1. To tackle this challenge, recent efforts have concentrated on designing memory systems tailored to embedding operation (kwon2019tensordimm, ), (ke2020recnmp, ), (park2021trim, ), (kal2021space, ), (ke2021near, ), (asgari2021fafnir, ). Approaches like TensorDIMM (kwon2019tensordimm, ) and RecNMP (ke2020recnmp, ) have explored the effectiveness of incorporating near-memory-processing (NMP) to alleviate the memory-bound nature by accelerating embedding operations directly within off-chip memory. These architectural solutions expedite the gather-and-reduce (GNR) process of embedding operations by deploying processing units in each rank of Dual Inline Memory Module (DIMM). This approach effectively harnesses the internal memory bandwidth, further amplified by the number of ranks in the memory channel. TRiM (park2021trim, ) and SPACE (kal2021space, ) followed a similar NMP structure while introducing other modifications to memory architecture. TRiM employed bank group-level processing-in-memory (PIM) within the DRAM device, while SPACE utilized High Bandwidth Memory (HBM) as a Dynamic Random Access Memory (DRAM) cache.
@@
Figure 1. Model size and bandwidth growth of DLRM. (Reproduction from RecShard (sethi2022recshard, ))
@@
On the other hand, personalized recommendation models are currently expanding in size. According to RecShard (sethi2022recshard, ), the memory capacity requirement of the Deep Learning Recommendation Model (DLRM) has been enlarged 16 times from 2017 to 2021, as shown in Figure 1. The model’s growth is attributed to the fact that more sparse features attribute to better model quality (sethi2022recshard, ), (mudigere2022software, ). However, The model operating on the inference server (zhao2022understanding, ), (ardestani2022supporting, ), (ke2022hercules, ), (firoozshahian2023mtia, ) is significantly smaller in scale compared to the training model (sethi2022recshard, ), (mudigere2022software, ). Therefore, a significant gap exists between the training model size and the inference model size. Serving the inference with a multi-node server or introducing Solid-State-Drivers (SSDs) could be a possible solution. Still, it comes at the cost of synchronization and vulnerability to failures for the multi-node server case, and the SSDs have a negative impact on execution time (ardestani2022supporting, ), (zhao2019aibox, ). Therefore, exploring an alternative approach is essential to optimize model inference without encountering these drawbacks. Prior works took algorithmic approaches to reduce the embedding table size (shi2020compositional, ), (ginart2021mixed, ), (yin2021tt, ), (lyu2022optembed, ). While these techniques have successfully reduced the size of the embedding table, their work has been done at the cost of imposition of a much heavier burden on the memory bandwidth or inefficient memory capacity management. A representative method is compositional embedding (shi2020compositional, ), which employs a simple double-hashing technique to transform the original embedding table into a smaller quotient (Q) table and remainder (R) table. Nevertheless, this method necessitates double memory access, amplifying the memory-bound characteristics.
@@
We propose HEAM, a specialized memory architecture integrated with PIM technology, designed to accelerate the inference process of compositional embedding. HEAM introduces a three-tiered memory hierarchy consisting of DIMM, HBM, and a lookup table (LUT). HBM is incorporated with base-die PIM (bd-PIM), while LUT serves as the storage space within bank-group PIM (bg-PIM), enhancing memory parallelism. The GnR processing in PIM can further reduce memory bandwidth requirements. This design is based on the following observations. Similar to original embedding, Q table exhibits high temporal locality characteristic within a small portion of entire embedding vector. Therefore, HEAM leverages HBM and DIMM to exploit this temporal locality. Notably, the R table, in particular, has a very small size and demonstrates a remarkably high temporal locality. Hence, we utilize an LUT within bg-PIM to store the entire R table. By addressing the memory bandwidth challenges associated with compositional embedding, HEAM effectively reduces the size of the embedding tables while simultaneously satisfying the requirements for memory bandwidth. This capability enables the execution of large models on a single-node server at a high processing speed. The primary contributions of our work can be summarized as follows:
@@
  • We design a two-level PIM within HBM, specifically tailored for compositional embedding. This enables HEAM to expand into a three-tiered memory system, incorporating DIMM, HBM, and the LUT within PIM, thus effectively enhancing memory parallelism.
2. Background
2.1. Personalized Recommendation System
@@
Figure 2. DLRM architecture
@@
Challenges in Recommendation System Inference. The performance of the recommendation system is constrained by the memory-bound nature of the embedding operation. The irregular access patterns of embedding vectors, combined with the extensive size of the embedding table that surpasses the cache storage capacity on the host side, necessitate the frequent look-up of embedding vectors from the off-chip memory system. This frequent access to main memory leads to higher latency for the embedding operation than that of the bottom-mlp (kal2021space, ), classifying the model as memory-bound. Recently, recommendation models have been trending towards incorporating increasingly large-sized embedding tables, with their scale growing year by year (sethi2022recshard, ). Nowadays, industry-scale DLRM implementations require substantial memory resources, often reaching sizes of up to tens of terabytes, a significant portion of which is dedicated to housing embedding tables (ardestani2022supporting, ), (zhao2022understanding, ). Given that the quality of the model is closely linked to the size of these embedding tables, there is a strong incentive to utilize as large embedding tables as possible in production. However, the practical utilization of such large models is constrained by physical limitations, as the memory size of a single-node server is restricted by the available slots within the server.
2.2. Algorithms for Reducing Embedding Data Size
@@
Algorithms Overview. Table 1 presents a comparison between aspects of algorithmic approaches for reducing embedding table size, including compression rates (Compression), accuracy relative to the original model (AUC), and operational characteristics such as bandwidth bound (BW bound) and memory friendliness (Mem. friendly). These methods could be categorized into three groups: Parameter Sharing (PS), Mixed Dimension (MD) and Generative Embeddings (GEN). The PS group utilizes hashing methods that map certain subsets of embedding vectors to the same vector. Conversely, MD group focuses on eliminating redundancy within the embedding dimensions, where embeddings that have minimal impact on the output are resized to a smaller dimension. Instead of storing embeddings as data, GEN group generates embedding from input indices using compute-intensive operations such as MLP. The PS group offers advantages regarding memory efficiency as its methods produce vectors with uniform dimensions as the algorithm’s output, promoting efficient utilization of memory capacity and accommodating memory system bursts effectively. Another benefit is the relatively low computational complexity of hash functions. However, the drawback is that the accuracy lags behind the baseline due to the semantic overlap among non-relevant embeddings. Furthermore, the degree of compression achieved is relatively modest. The MD and GEN groups attain a high compression rate and better model accuracy than the baseline. In the MD group, however, efficient use of memory resources is relatively hard since the output vectors vary in their dimension. Finally, the GEN group transforms the target model into the compute-bound application, which is far from the memory perspective, in contrast to other memory-bound groups. The MD and GEN groups achieve a substantial compression rate while delivering improved model accuracy compared to the baseline. In the MD group, however, efficient utilization of memory resources is relatively challenging because the output vectors exhibit varying dimensions. On the other hand, the GEN group transforms the target model into a compute-bound application, diverging significantly from the memory-oriented perspective of the other groups, which are memory-bound applications.
Table 1. Comparison of embedding table reduction schemes.
@@
PS Hashing (weinberger2009feature, ) Low Much Lower Memory
QR (shi2020compositional, ) Moderate Lower Memory
MD MDE (ginart2021mixed, ) Moderate Better Memory
OptEmbed (lyu2022optembed, ) High Better Memory
GEN TT-Rec (yin2021tt, ) High Better Compute
DHE (kang2021learning, ) High Better Compute
@@
Compositonal Embedding. Compositional embedding employs two types of hash functions, effectively partitioning the original embedding table into two smaller tables, as depicted in Figure 3 (a). These hash functions include the quotient operation and remainder operation. Embeddings that yield the same quotient result are mapped to corresponding vectors in the table generated by the quotient function (Q table). In contrast, embeddings with identical remainder results are directed to the same vector within the other hash table (R table). The total table size is typically reduced by a factor approximately equal to the inverse of the hash collision, with the R table size being usually negligible compared to that of the Q table. The R table size corresponds to the hash collision, whereas the Q table size is equivalent to the original table size divided by the hash collision. To process the embedding operation, it is necessary to reconstruct the original embedding. This is accomplished by performing straightforward mathematical operations on the vectors extracted from each table. The overall scheme is illustrated in Figure 3 (b).
@@
Figure 3. Compositional embedding overview. (a) Double hashing process (b) Embedding vector reconstruction process
2.3. NMP/PIM Designs for Recommendation System
@@
Recent studies have introduced NMP units within the DRAM architecture to meet the memory bandwidth requirements of recommendation systems. Specifically, these approaches have incorporated multiple DIMMs with NMP units situated within their DIMM/Rank configurations, effectively exploiting DIMM/Rank-level parallelism (kwon2019tensordimm, ), (ke2020recnmp, ) to achieve increased throughput. TRiM (park2021trim, ) focused on the tree topological structure of the DRAM’s data path, further enhancing parallelism by integrating PIM units within each bank group. In pursuit of additional speedup, the aforementioned designs have leveraged the long-tail distribution of embeddings, where a small subset of embeddings is frequently reused. This has been accomplished by incorporating a cache within the buffer chip (ke2020recnmp, ) or implementing load-balancing techniques with the assistance of the memory controller (park2021trim, ).
@@
While it is possible to increase the throughput of the embedding operation by incorporating more DIMMs with NMP designs from previous approaches, this may lead to inefficiencies in power consumption (kal2021space, ). To address this challenge, SPACE (kal2021space, ) introduced a heterogeneous memory system that combines HBM and DIMMs, with HBM functioning as a DRAM cache. Additionally, SPACE exploited the reduced locality of embedding vectors and the long-tail distribution of embedding vectors to achieve significantly higher throughput. In summary, achieving high performance in recommendation systems relies on core techniques that involve exploiting parallelism within the main memory and capitalizing on the long-tail distribution of embedding vectors.
3. Analysis on Compositional Embedding
Figure 4. Compositional Embedding Analysis (a) Execution time comparison between original DLRM and compositional-embedding-applied DLRM (b) Verifying temporal locality by increasing cache size from 1MB to 4MB with vector length of 64B (c) Spatial locality by increasing cache line size from 64B to 512B where cache size is fixed to 1MB with vector length of 64B
@@ 3. Analysis on Compositional Embedding
As shown in Section 2.2, compositional embedding exhibits several advantages over other previously proposed schemes concerning the memory system. While alternative methods may yield better model prediction quality and higher compression rates, the nature of the output data poses challenges for the memory system when supporting the embedding operation. Consequently, we have selected compositional embedding as the target algorithm. In this section, we offer an in-depth analysis of the characteristics of compositional embedding.
3.1. Impact on DLRM Throughput

### 14 Mar 2024 v3

@@ Abstract
In today’s data centers, personalized recommendation systems face challenges of large memory capacity requirements and high memory bandwidth demand, especially for embedding operations. Previous approaches have relied on DIMM-based near-memory processing techniques or introduced 3D-stacked DRAM to address memory-bound issues. However, these solutions fall short when dealing with the expanding size of personalized recommendation systems. Recommendation models have grown to sizes exceeding tens of terabytes, making them challenging to run efficiently on traditional single-node inference servers. Although various algorithmic methods have been proposed to reduce embedding table capacity, they often result in increased memory access or inefficient utilization of memory resources.
@@
This paper introduces HEAM, a heterogeneous memory architecture that integrates 3D-stacked DRAM with DIMM to accelerate recommendation systems in which compositional embedding is utilized—a technique to reduce the size of embedding tables. The architecture is organized into a three-tier memory hierarchy consisting of conventional DIMM, 3D-stacked DRAM with a base die-level Processing-In-Memory (PIM), and lookup tables inside bank group-level PIM. This setup is specifically designed to accommodate the unique aspects of compositional embedding, such as temporal locality and embedding table capacity. This design effectively reduces bank access, improves access efficiency, and enhances overall throughput, resulting in 6.2×6.2$\times$6.2 × speedup and 58.9% energy savings compared to the baseline.
Index Terms:
Processing in memory, recommendation system, compositional embedding, heterogeneous memory, 3D-stacked DRAM.
I Introduction
@@
Personalized recommendation system serves as an essential technology of various companies such as Meta [1], YouTube [2] and Netflix [3]. Recently, deep learning-based recommendation systems have emerged as a pivotal technology for improving the accuracy of predicting user-preferred content, thereby contributing to enhanced profitability for these companies. These advanced recommendation systems currently contribute substantially to the computational workload of the data center. Recent reports indicate that approximately 80% of data center resources are dedicated to the inference process of recommendation systems [4]. Recommendation models use dense features for user information and sparse features for item specifics through embedding operations. Following feature interactions, multi-layer perceptron (MLP) layers generate click-through-rate (CTR) predictions, estimating the likelihood of user clicks on recommended items.
@@
In large-scale data centers that extensively utilize embeddings, the embedding operation often becomes the primary factor influencing overall system performance. This stems mainly from the operation’s tendency to exhibit sparse and irregular memory access patterns. As a result, these memory-bound operations place significant constraints on traditional data center infrastructures [5], and this challenge continues to grow with the scale, as depicted in Figure 1. To tackle this challenge, recent efforts have concentrated on designing memory systems tailored to embedding operation [6], [7], [8], [9], [10], [11]. Approaches like TensorDIMM [6] and RecNMP [7] have explored the effectiveness of incorporating near-memory-processing (NMP) to alleviate the memory-bound nature by accelerating embedding operations directly within off-chip memory. These architectural solutions expedite the gather-and-reduce (GNR) process of embedding operations by deploying processing units in each rank of Dual Inline Memory Module (DIMM). This approach effectively harnesses the internal memory bandwidth, further amplified by the number of ranks in the memory channel. TRiM [8] and SPACE [9] followed a similar NMP structure while introducing other modifications to memory architecture. TRiM employed bank group-level processing-in-memory (PIM) within the DRAM device, while SPACE utilized High Bandwidth Memory (HBM) [12] as a Dynamic Random Access Memory (DRAM) cache.
@@
Figure 1: Model size and bandwidth growth of DLRM. (Reproduction from RecShard [13])
@@
On the other hand, personalized recommendation models are currently expanding in its size. According to RecShard [13], the memory capacity requirement of the Deep Learning Recommendation Model (DLRM) has been enlarged 16 times from 2017 to 2021, as shown in Figure 1. The model’s growth is attributed to the fact that incorporating more items attributes to better model quality [13], [14]. However, the model operating on the inference server [5], [15], [16], [17] is significantly smaller in scale compared to the training model [13], [14]. Therefore, a significant gap exists between the training model size and the inference model size. Serving the inference with a multi-node server or introducing Solid-State-Drivers (SSDs) could be a possible solution. However, it comes at the cost of synchronization and vulnerability to failures for the multi-node server case, and the SSDs have a negative impact on execution time [15], [18]. Therefore, exploring an alternative approach is essential to optimize model inference without encountering these drawbacks. Prior works took algorithmic approaches to reduce the embedding table size [19], [20], [21], [22]. While these techniques have successfully reduced the size of the embedding table, their work has been done at the cost of imposition of a much heavier burden on the memory bandwidth or inefficient memory capacity management. A representative method is compositional embedding [19], which employs a simple double-hashing technique to transform the original embedding table into a smaller quotient (Q) table and remainder (R) table. Nevertheless, this method necessitates double memory access, amplifying the memory-bound characteristics.
@@
We propose HEAM, a specialized memory architecture integrated with PIM technology, designed to accelerate the inference process of compositional embedding. HEAM introduces a three-tiered memory hierarchy consisting of DIMM, HBM, and a lookup table (LUT). HBM is incorporated with base-die PIM (bd-PIM), and bank-group PIM (bg-PIM), and LUT that serves as an additional storage space within bg-PIM. This design is based on the following observations. Similar to the original embedding, the Q table exhibits high temporal locality characteristics within a small portion of the entire embedding vector. Therefore, HEAM leverages HBM and DIMM to exploit this temporal locality, storing high temporal locality vectors to the HBM and others to the DIMM. Notably, the R table, in particular, has a very small size and demonstrates a remarkably high temporal locality. Hence, we utilize an LUT within bg-PIM to store the entire R table. By addressing the memory bandwidth challenges associated with compositional embedding, HEAM effectively reduces the size of the embedding tables while simultaneously satisfying the requirements for memory bandwidth. This capability enables the execution of large models on a single-node server at a high processing speed. The primary contributions of our work can be summarized as follows:
@@
  • We design a two-level PIM within HBM, specifically tailored for compositional embedding. This enables HEAM to expand into a three-tiered memory system, incorporating DIMM, HBM, and the LUT within PIM, thus effectively enhancing memory parallelism. Using LUT facilitates the additional speedup of Processing-In-Memory (PIM) architectures.
II Background
II-A Personalized Recommendation System
@@
Figure 2: DLRM architecture
@@
Challenges in Recommendation System Inference. The performance of the recommendation system is constrained by the memory-bound nature of the embedding operation. The irregular access patterns of embedding vectors, combined with the extensive size of the embedding table that surpasses the cache storage capacity on the host side, necessitate the frequent look-up of embedding vectors from the off-chip memory system. This frequent access to main memory leads to higher latency for the embedding operation than that of the bottom-mlp [9], classifying the model as memory-bound. Recently, recommendation models have been trending towards incorporating increasingly large-sized embedding tables, with their scale growing year by year [13]. Nowadays, industry-scale DLRM implementations require substantial memory resources, often reaching sizes of up to tens of terabytes, a significant portion of which is dedicated to housing embedding tables [15], [5]. Given that the quality of the model is closely linked to the size of these embedding tables, there is a strong incentive to utilize as large embedding tables as possible in production. However, the practical utilization of such large models is constrained by physical limitations, as the memory size of a single-node server is restricted by the available slots within the server.
II-B Algorithms for Reducing Embedding Data Size
@@
Algorithms Overview. Table I presents a comparison between aspects of algorithmic approaches for reducing embedding table size, including compression rates (Compression), accuracy relative to the original model (AUC), and operational characteristics such as bandwidth bound (BW bound) and memory friendliness (Mem. friendly). These methods could be categorized into three groups: Parameter Sharing (PS), Mixed Dimension (MD) and Generative Embeddings (GEN). The PS group utilizes hashing methods that map consecutive embedding vectors to the same vector. Conversely, MD group focuses on eliminating redundancy within the embedding dimensions, where embedding vectors that have minimal impact on the output are resized to a smaller dimension. The GEN group generates embedding from input indices using compute-intensive operations such as MLP, instead of storing embedding vectors as data. The PS group offers advantages regarding memory efficiency as its methods produce vectors with uniform dimensions as the algorithm’s output, promoting efficient utilization of memory capacity and accommodating memory system bursts effectively. However, the drawback is that the accuracy lags behind the baseline due to the semantic overlap among non-relevant embedding vectors. Furthermore, the degree of compression achieved is relatively modest. The MD and GEN groups attain a high compression rate and better model accuracy than the baseline. In the MD group, however, efficient use of memory resources is relatively hard since the output vectors vary in their dimension. Finally, the GEN group transforms the target model into the compute-bound application, which is far from the memory perspective, in contrast to other memory-bound groups.
TABLE I: Comparison of embedding table reduction schemes.
@@
PS Hashing [23] Low Much Lower Memory
QR [19] Moderate Lower Memory
MD MDE [20] Moderate Better Memory
OptEmbed [22] High Better Memory
GEN TT-Rec [21] High Better Compute
DHE [24] High Better Compute
@@
Compositonal Embedding. Compositional embedding employs two types of hash functions, effectively partitioning each of the original embedding tables into two smaller tables, as depicted in Figure 3 (a). These hash functions consist of quotient and remainder operations, where the original embedding vector index is used as an input. Embedding vectors that yield the same quotient value, that are consecutive embedding vectors, are mapped to corresponding vectors in the table generated by the quotient function (Q table). On the other hand, embedding vectors with identical remainder results, which departed from each other with the distance of hash collision value, are directed to the same vector within the other hash table (R table). To process the embedding operation, it is necessary to reconstruct the original embedding from Q table and R table. This is accomplished by performing straightforward mathematical operations on the vectors extracted from each table. The overall scheme is illustrated in Figure 3 (b). Note that the user may select reconstruction operation among addition, multiplication and concatenation[19]. The total table size is reduced by a factor approximately equal to the inverse of the hash collision. The R table size is usually negligible compared to the Q table, as its size corresponds to the hash collision. In contrast, the Q table size is equivalent to the original table size divided by the hash collision.
@@
Figure 3: Compositional embedding overview. (a) Double hashing process (b) Embedding vector reconstruction process
II-C NMP/PIM Designs for Recommendation System
@@
Recent studies have introduced NMP units within the DRAM architecture to meet the memory bandwidth requirements of recommendation systems. Specifically, these approaches have incorporated multiple DIMMs with NMP units situated within their DIMM/Rank configurations, effectively exploiting DIMM/Rank-level parallelism [6], [7] to achieve increased throughput. TRiM [8] focused on the tree topological structure of the DRAM’s data path, further enhancing parallelism by integrating PIM units within each bank group. In these previous designs, NMP/PIM operations are not executed in a synchronized style as embedding vectors are accessed randomly. However, as each vector read requires more than one DRAM read access cycle (tRAS), a new NMP/PIM operation could be successfully transferred and initiated before the previous instruction in another NMP/PIM unit completes, resulting in parallel executions. In pursuit of additional speedup, the aforementioned designs have leveraged the long-tail distribution of embedding vectors, where a small subset of embedding vectors is frequently reused. This has been accomplished by incorporating a cache within the buffer chip [7] or implementing load-balancing techniques with the assistance of the memory controller [8].
@@
While it is possible to increase the throughput of the embedding operation by incorporating more DIMMs with NMP designs from previous approaches, this may lead to inefficiencies in power consumption [9]. To address this challenge, SPACE [9] introduced a heterogeneous memory system that combines HBM and DIMMs, with HBM functioning as a DRAM cache. Additionally, SPACE exploited the reduction locality of embedding vectors and the long-tail distribution of embedding vectors to achieve significantly higher throughput. In summary, achieving high performance in recommendation systems relies on core techniques that involve exploiting parallelism within the main memory and capitalizing on the long-tail distribution of embedding vectors.
III Analysis on Compositional Embedding
@@
As shown in Section II-B, compositional embedding exhibits several advantages over other previously proposed schemes concerning the memory system. While alternative methods may yield better model prediction quality and higher compression rates, the nature of the output data poses challenges for the memory system when supporting the embedding operation efficiently. Consequently, we have selected compositional embedding as the target algorithm. In this section, we offer an in-depth analysis of the characteristics of compositional embedding.
Figure 4: Compositional Embedding Analysis (a) Execution time comparison between original DLRM and compositional-embedding-applied DLRM (b) Verifying temporal locality by increasing cache size from 1MB to 4MB with vector length of 64B (c) Spatial locality by increasing cache line size from 64B to 512B where cache size is fixed to 1MB with vector length of 64B
III-A Impact on DLRM Throughput